# FEDONet: Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v3

@@ 3.5 Allen–Cahn Equation
We further evaluate the performance of the proposed FEDONet in modeling the one-dimensional Allen–Cahn nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
@@
where $u(x,t)$ is the phase field variable and $\epsilon=10^{-4}$ denotes the diffusion coefficient. Periodic boundary conditions are enforced on both $u$ and its spatial derivative $\partial u/\partial x$, consistent with the physical assumption of a closed domain.
To construct a challenging dataset with rich spatial features, we generate 10,000 solution trajectories using an explicit Euler solver. The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$. Initial conditions are defined as:
@@
where $a_{k},b_{k}\sim\mathcal{U}(0,1)$ are independent samples from the uniform distribution. This construction introduces multiscale structure and localized steep gradients into the initial data, posing a nontrivial test for surrogate learning.
Each training sample comprises a pair $(s(x),u(x,t))$, where $s(x)\in\mathbb{R}^{200\times 1}$ denotes the initial profile, and $u(x,t)\in\mathbb{R}^{200\times 200}$ is the full spatiotemporal solution. The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
Figure 13: Solving a parametric Allen Cahn equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Both models were evaluated on $128$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $10.61\%$ for the DeepONet and $5.50\%$ for the FEDONet. Figure 13 shows qualitative predictions for a representative test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The pointwise absolute error plot in Figure 14 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 14: Energy spectrum and pointwise absolute error across space and time for a representative sample for Allen-Cahn Equation.
Figure 14 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen-Cahn system.
3.6 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
@@ 3.6 Kuramoto–Sivashinsky Equation
where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series:
@@
which introduces low-frequency coherent structures that quickly evolve into chaotic waveforms under the KS dynamics.
We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver. The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the proposed FEDONet models. This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
Figure 15: Solving a parametric KS equation: Median performing test-sample. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 15 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 16, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
@@
Figure 17 displays the worst-case test result. While the FEDONet prediction exhibits growing deviations from the reference field, it still retains discernible wave structures and phase coherence, with an error of $58.02\%$. In comparison, the vanilla DeepONet fails entirely, producing numerical artifacts, amplitude distortion, and full trajectory collapse. This robustness under extreme conditions emphasizes the value of frequency-enhanced representations in learning dynamical system operators.
Figure 18 plots the training loss curves for both models. FEDONet converges faster and exhibits more stable optimization behavior, attributable to the improved conditioning of the input space via randomized Fourier features. This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
Figure 18: KS Training convergence comparison. FEDONet achieves faster and more stable optimization.
Figure 19: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
Finally, Figure 19 summarizes performance statistics across 512 randomly selected test cases. FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
In totality, the Kuramoto-Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
@@ 4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations. By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $L^{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $L^{2}$ error (%) for each benchmark.
Dataset DeepONet FEDONet
2-D Poisson (elliptic) 1.41 1.09
1-D Burgers (conv./diff.) 6.43 4.86
Lorenz-63 (chaotic ODE) 2.03 0.46
Allen-Cahn (phase–field) 10.61 5.50
Kuramoto-Sivashinsky (spatio-chaos) 75.11 20.62
Eikonal (Hamilton–Jacobi) 2.15 1.12
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity. Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder-decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework via Bayesian formulations or PINO style residual losses may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
Declaration of competing interest
@@ Acknowledgements
Data supporting the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
@@
To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is:
@@
We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
@@
Therefore, the unscaled covariance matrix for this block is approximately:
@@
Scaling by the $\sqrt{2}$ factor yields:
@@
Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
@@
Appendix B Theoretical Justification for Fourier Embedding Superset Property
In this appendix, we formally justify the claim stated in Section 2.1 that the hypothesis space induced by Fourier embeddings strictly contains the vanilla hypothesis space of coordinate-input networks. We present this as a lemma and provide constructive proof using classical function approximation theory.
Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z\in\mathbb{R}^{M\times d},\;Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (39)
Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (40)
We divide the proof into two parts.
Part I: Inclusion. Let $f\in\mathcal{H}_{\text{vanilla}}$ be a function representable by a neural network of the form:
$f(\zeta)=\sum_{j=1}^{N}a_{j}\,\sigma(w_{j}^{\top}\zeta+b_{j}).$ (41)
Since each $\zeta_{i}$ (coordinate function) can be approximated by a Fourier series on a compact domain $K$, and the Fourier embedding $\phi(\zeta)$ contains these sinusoidal components, it follows that a function of $\zeta$ can be emulated by a function of $\phi(\zeta)$. In particular, for every $f\in\mathcal{H}_{\text{vanilla}}$ and $\varepsilon>0$, there exists $\tilde{f}\in\mathcal{H}_{\phi}$ such that:
$\sup_{\zeta\in K}|f(\zeta)-\tilde{f}(\zeta)|<\varepsilon.$ (42)
Therefore, $\mathcal{H}_{\text{vanilla}}\subseteq\mathcal{H}_{\phi}$.
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [51], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $Z$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi Z_{j}\zeta)\approx\sin(100\pi\zeta),$
making $f$ easily representable in $\mathcal{H}_{\phi}$.
Since $f(\zeta)$ cannot be efficiently approximated in $\mathcal{H}_{\text{vanilla}}$ (requiring significantly more parameters or depth), but is trivially accessible in $\mathcal{H}_{\phi}$, we conclude that the inclusion is strict:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (43)
This result formalizes the intuition that Fourier embeddings expand the expressivity of the network and can be interpreted as a form of spectral preconditioning for operator learning models such as DeepONets.
Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
The use of Fourier embeddings in the trunk network induces a natural spectral decomposition in the DeepONet output. Specifically, the composition of the trunk and branch networks can be interpreted as a data-driven spectral projection:
$\mathcal{G}(u)(\zeta)\approx\sum_{k=1}^{p}\langle u,\varphi_{k}\rangle\,\psi_{k}(\zeta),$ (44)
where $\{\varphi_{k}\}_{k=1}^{p}$ are data-dependent basis functions learned from the branch network, and $\{\psi_{k}(\zeta)\}_{k=1}^{p}$ are basis functions defined by the trunk network, shaped by the sinusoidal feature embedding $\phi(\zeta)$.
When the trunk network is fed sinusoidal embeddings of the form
$\phi(\zeta)=[\cos(2\pi\omega_{1}^{\top}\zeta),\dots,\cos(2\pi\omega_{M}^{\top}\zeta),\sin(2\pi\omega_{1}^{\top}\zeta),\dots,\sin(2\pi\omega_{M}^{\top}\zeta)],$
the basis functions $\psi_{k}(\zeta)$ are learned combinations of trigonometric modes. This allows the model to represent oscillatory and localized patterns that are characteristic of multiscale physical systems.
The representation in Equation (44) mirrors classical Galerkin methods, where the solution is projected onto a known orthonormal basis:
$u(x,t)\approx\sum_{n=1}^{N}a_{n}(t)\,\phi_{n}(x),$ (45)
with the coefficients $a_{n}(t)$ governed by PDE dynamics. In DeepONet, both the basis functions $\psi_{k}$ and the projection coefficients $\langle u,\varphi_{k}\rangle$ are learned directly from data, enabling generalization beyond handcrafted bases.
Unlike classical global bases such as Fourier or Chebyshev functions, the neural trunk network can generate localized, adaptive basis functions. This enhances the model’s ability to approximate functions with sharp transitions, discontinuities, or compactly supported features, which are common in real-world PDE systems. We may formalize the DeepONet architecture as a composition of a projection and synthesis operator:
$\Pi_{\varphi}(u):=[\langle u,\varphi_{1}\rangle,\dots,\langle u,\varphi_{p}\rangle]^{\top},\quad\mathcal{S}_{\psi}(\zeta):=[\psi_{1}(\zeta),\dots,\psi_{p}(\zeta)]^{\top},$
yielding the output:
$\mathcal{G}(u)(\zeta)\approx\mathcal{S}_{\psi}(\zeta)^{\top}\Pi_{\varphi}(u),$
which resembles a spectral synthesis operation of the form:
$f(x)\approx\sum_{k}\hat{f}_{k}\phi_{k}(x).$
This interpretation leads to several insights:
  1. 1.
    Fourier-Embedded DeepONets naturally align with the dominant spectral modes of the target operator, facilitating more efficient approximation of high-frequency content.
  2. 2.
    The learned basis can adapt to complex patterns in the data, unlike fixed basis expansions.
  3. 3.
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov-Galerkin and spectral methods, but operates in a fully data-driven regime.
In summary, Fourier Embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators. This perspective bridges classical PDE approximation theory with modern deep learning, offering both interpretability and flexibility in learning complex function-to-function mappings.

### v4

@@ 3.4 Allen-Cahn Equation
We further evaluate the performance of FEDONet in modeling the 1D Allen-Cahn nonlinear reaction-diffusion system, widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by,
@@
where $u(x,t)$ is the phase field variable and $\epsilon=10^{-4}$ denotes the diffusion coefficient. Periodic boundary conditions are enforced on both $u$ and its spatial derivative $\partial u/\partial x$, consistent with the physical assumption of a closed domain. To construct a challenging dataset with rich spatial features, we generate 10,000 solution trajectories using an explicit Euler solver. The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$. Initial conditions are defined as,
@@
where $a_{k},b_{k}\sim\mathcal{U}(0,1)$ are independent samples from the uniform distribution. This construction introduces multiscale structure and localized steep gradients into the initial data, posing a nontrivial test for surrogate learning. Each training sample comprises a pair $(s(x),u(x,t))$, where $s(x)\in\mathbb{R}^{200\times 1}$ denotes the initial profile, and $u(x,t)\in\mathbb{R}^{200\times 200}$ is the full spatiotemporal solution. The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
Figure 12: Solving a parametric Allen-Cahn equation: Comparison of predicted spatio-temporal solution fields for a test input with additive noise. The corresponding relative $L^{2}$ errors are $7.64\%$ for DeepONet and $3.16\%$ for FEDONet.
Both models were evaluated on $250$ randomly selected, unseen, noise-free test samples, yielding average relative $L^{2}$ errors of $10.98\%$ for DeepONet and $5.85\%$ for FEDONet. Figure 12 presents qualitative predictions for a representative test case with $10\%$ additive noise in the input. While DeepONet captures the overall structure of the solution, it exhibits noticeable smoothing and loss of sharpness near interface regions. In contrast, FEDONet produces significantly sharper reconstructions, accurately resolving interface zones and phase transitions with minimal blurring.
Figure 13: Spectral and Error analysis for Allen-Cahn equation: Comparison of the energy spectrum and pointwise absolute error fields for a representative noise-free test sample.
The pointwise absolute error profiles for a representative noise-free test case, shown in Figure 13, highlight the localized accuracy of FEDONet in resolving steep interface regions. DeepONet exhibits significant error accumulation in regions with sharp spatial gradients, particularly near evolving interfaces, reflecting its tendency to under-resolve high-frequency components. In contrast, FEDONet maintains consistently low error across the domain, with only minor localized deviations, indicating a superior ability to capture sharp transitions and preserve interface structure. This behavior is further corroborated by the spectral analysis in Figure 13, where the energy spectra of the predicted and true solutions are compared. FEDONet closely tracks the ground truth spectrum, whereas DeepONet exhibits noticeable spectral attenuation for $k>20$.
3.5 Kuramoto-Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by,
@@ 3.5 Kuramoto-Sivashinsky Equation
where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains, respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series,
@@
which introduces low-frequency coherent structures that quickly evolve into chaotic waveforms under the KS dynamics. We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver. The resulting spatial and temporal grids are subsequently normalized to the unit domain $0,1]\times[0,1]$ for operator learning. The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the FEDONet. This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
Figure 14: Solving a parametric KS equation: Median performing test-sample. The corresponding relative $L^{2}$ errors are 89.63% for DeepONet and 16.36% for FEDONet.
Figure 14 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse. This degradation is reflected in a high average relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet for 512 unseen noise-free test samples.
@@
In Figure 15, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift.
Figure 16: Robustness to input noise for the KS equation: Qualitative comparison of predicted spatiotemporal fields under increasing noise levels in the initial condition.
The quantitative impact of input noise is summarized in Fig. 16. DeepONet exhibits consistently high error even in the noise-free setting and shows only marginal variation with increasing noise, indicating a fundamental inability to capture the underlying operator. In contrast, FEDONet achieves substantially lower error across all noise levels, although a gradual increase is observed as noise increases. Notably, even at high noise levels (up to $30\%$), FEDONet significantly outperforms DeepONet, demonstrating superior robustness and stability in chaotic regimes.
@@ 4 Summary
In this work, we introduced the Fourier-Embedded DeepONet (FEDONet), a principled extension of the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations. By applying randomized Fourier feature embeddings to the trunk coordinates, FEDONet performs a spectral lifting of the input space that can be interpreted as a form of spectral preconditioning, improving the conditioning of the learning problem, expanding the effective hypothesis space, and mitigating the spectral bias of standard multilayer perceptrons. This enables efficient representation of multiscale and oscillatory solution structures while preserving the original branch-trunk formulation and incurring negligible computational overhead. Extensive evaluations across a diverse suite of canonical PDE benchmarks demonstrate that FEDONet consistently outperforms the vanilla DeepONet across elliptic, parabolic, hyperbolic, and chaotic regimes, with particularly significant gains in systems exhibiting broadband frequency content such as the Kuramoto-Sivashinsky equation. A comprehensive summary of the results is provided in Table 1. Beyond pointwise accuracy, FEDONet exhibits superior spectral fidelity, accurately capturing inertial and dissipative ranges without artificial high-frequency energy accumulation, and demonstrates enhanced robustness to input noise as well as improved sample efficiency in low-data regimes. Unlike global spectral operator methods such as Fourier Neural Operators, FEDONet retains the locality and modularity of the DeepONet framework while introducing a principled spectral inductive bias, establishing it as a simple, scalable, and architecture-agnostic enhancement for high-fidelity operator learning in scientific machine learning.
Table 1: Mean relative $L^{2}$ error (%) of FEDONet and DeepONet.
Dataset DeepONet FEDONet Improvement
1-D Burgers (convection-diffusion) 6.43 4.86 24.4%
2-D Poisson (elliptic) 1.41 1.09 22.7%
Eikonal (Hamilton-Jacobi) 2.15 1.12 47.9%
Allen-Cahn (phase-field) 10.98 5.85 46.72%
Kuramoto-Sivashinsky (spatio-temporal chaos) 89.63 16.36 81.75%
5 Conclusion and Future Work
While FEDONet demonstrates strong performance across a range of operator learning tasks, several directions remain for further development, particularly toward real-world applicability. A natural extension is to move beyond fixed sinusoidal embeddings by exploring alternative spectral or localized basis functions, such as wavelets or orthogonal polynomials, to improve spatial localization and boundary adaptivity. Integrating the spectral embedding strategy with structured architectures, including graph-based trunks or multiscale encoder-decoder frameworks, also offers a pathway toward handling complex geometries and irregular domains. Another promising direction is the introduction of adaptive or learnable spectral embeddings, which could better align with the intrinsic frequency content of the underlying operator and improve data efficiency. In addition, incorporating uncertainty quantification and physics-based constraints will be essential for robust deployment in practical settings.
Although this study focuses on canonical PDE benchmarks, FEDONet is motivated by challenges in high-dimensional systems such as turbulent flows and geophysical dynamics, where multiscale structure, noise sensitivity, and spectral complexity are dominant. The Fourier-embedded trunk provides a lightweight mechanism for introducing spectral inductive bias, making the framework well-suited for applications in surrogate modeling and real-time prediction. Key challenges remain, including scalability to high-resolution domains, handling complex geometries, and integration with data assimilation or hybrid physics-informed frameworks. Addressing these will be critical for extending operator learning methods from controlled benchmarks to realistic scientific applications.
Declaration of Competing Interests
@@ Acknowledgements
The datasets and code used to generate the results presented in this study are publicly available and can be accessed through the following GitHub repository: https://github.com/as26101999/Fourier-Embedded-DeepONets
@@ Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution,
@@
To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is,
@@
We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain,
@@
Therefore, the unscaled covariance matrix for this block is approximately,
@@
Scaling by the $\sqrt{2}$ factor yields,
@@
Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies,
@@
Appendix B Additional Results
B.1 Burgers’ Equation
Figure 17: Robustness to input noise for Burgers’ equation: Mean relative $L^{2}$ error as a function of input noise level. FEDONet consistently outperforms DeepONet and exhibits slower error degradation.
B.2 2D Poisson Equation
Figure 18: Sample efficiency for the 2D Poisson equation: Mean relative $L^{2}$ error as a function of the number of training samples.
B.3 Eikonal Equation
Figure 19: Robustness to input noise for the Eikonal equation: Mean relative $L^{2}$ error as a function of input noise level. FEDONet shows consistently lower error and slower degradation compared to DeepONet, indicating improved robustness.
B.4 Allen-Cahn Equation
Figure 20: Sample efficiency for the Allen-Cahn equation: Mean relative $L^{2}$ error as a function of the number of training samples.
Figure 21: Robustness to input noise for Allen-Cahn equation: Mean relative $L^{2}$ error as a function of input noise level.
B.5 Kuramoto-Sivashinsky Equation
Figure 23: Robustness to input noise for KS equation: Mean relative $L^{2}$ error as a function of input noise level.